# Reachability for multiagent control systems via Lyapunov functions

### v1

@@ 4.3. The case $\varphi(y)=\alpha$
Proof.
The proof of this result is similar to the proof of the left-to-right implication in Proposition 4.8. Fix $\mu_{0}\in\mathcal{P}_{2}(\mathbb{R}^{d})$, $T>0$. Due to the different rate function (4.4), in this case we consider the functional $g:\mathcal{P}_{2}(\mathbb{R}^{d})\to\mathbb{R}$ defined by $g(\nu):=\alpha\,T+\mathrm{V}(\nu),$ and $U_{g}$ as in (4.11). By the Dynamic Programming Principle (cf. [JMQ20, Corollary 2 and Proposition 3]), there exists $\mu\in\mathcal{A}_{[0,T]}(\mu_{0})$ such that $U_{g}(t,\mu_{t})=U_{g}(s,\mu_{s})$ for all $0\leq s\leq t\leq T$. By Lemma 4.12 andTheorem 3.8, we get
$\alpha\,T+\mathrm{V}(\mu_{T})=U_{g}(T,\mu_{T})=U_{g}(s,\mu_{s})\leq\mathrm{S}(s,\mu_{s})=\alpha\,s+\mathrm{V}(\mu_{s}),\quad\forall\,s\in[0,T].$
By the arbitrariness of $T>0$ and since the gluing of admissible trajectories is still an admissible trajectory, we conclude as follows. There exists $\mu\in\mathcal{A}_{[0,+\infty)}(\mu_{0})$ such that
[…]
in particular this holds for $s=0$, leading to $\mathrm{V}(\mu_{t})\leq\mathrm{V}(\mu_{0})-\alpha t$. Hence, since $\mathrm{V}\geq 0$, we deduce that (4.25) holds with $s,t\in\left[0,\frac{\mathrm{V}(\mu_{0})}{\alpha}\right]$, thus concluding the proof. ∎
Remark 4.14.
Notice that, differently from Proposition 4.8, the reverse implication in Proposition 4.13 does not hold in general. Indeed, by Lemma 4.11, proving that $\mathrm{V}$ is a Lyapunov function for $\mathrm{F}$ with rate $\varphi$ is equivalent to prove that $\mathrm{S}:[0,T]\times\mathcal{P}_{2}(\mathbb{R}^{d})\to[0,+\infty)$ is a partial-viscosity supersolution of (4.21) for some $T>0$ independent on the state measure $\mu_{0}$ where $\mathrm{S}$ is evaluated.
Corollary 4.15.
Assume Hypothesis 2.1. Let $\mathrm{V}:\mathcal{P}_{2}(\mathbb{R}^{d})\to[0,+\infty)$ be a Lyapunov function for $\mathrm{F}$ with rate $\varphi$. Then $\mathrm{epi}\,(\mathrm{V})$ is finite-time viable for the coupled dynamics
(4.26) $\begin{cases}\partial_{t}\mu_{t}+\mathrm{div}\,(v_{t}\,\mu_{t})=0,&t>0,\quad\text{with }v_{t}(x)\in\mathrm{F}(x,\mu_{t}),\\ y^{\prime}(t)=-\varphi(y(t)),&t>0;\end{cases}$
in particular, for any $(\mu_{0},y_{0})\in\mathrm{epi}\,(\mathrm{V})$ there exists a solution $(\mu,y)$ of (4.26) defined in the time interval $[0,T_{0}]$, with $T_{0}:=\frac{\mathrm{V}(\mu_{0})}{\alpha}$ and with initial condition $(\mu_{0},y_{0})$, i.e. $\mu\in\mathcal{A}_{[0,T_{0}]}(\mu_{0})$ and $y(t;y_{0})=-\alpha t+y_{0}$ for $t\in[0,T_{0}]$, such that
$\left(\mu_{t},y(t;y_{0})\right)\in\mathrm{epi}\,(\mathrm{V}),\quad\forall\,t\in[0,T_{0}].$
Theorem 4.16.
Assume Hypothesis 2.1. Let $\bar{\nu}\in\mathcal{P}_{2}(\mathbb{R}^{d})$ and assume there exists a Lyapunov function $\mathrm{V}$ for $\mathrm{F}$ with rate $\varphi$ such that $\mathrm{V}(\bar{\nu})=0$, $\mathrm{V}(\nu)>0$ for any $\nu\neq\bar{\nu}$. Then $\bar{\nu}$ is reachable; in particular, given $\mu_{0}\in\mathcal{P}_{2}(\mathbb{R}^{d})$ there exists $\mu\in\mathcal{A}_{[0,T]}(\mu_{0})$ such that $\mu_{T}=\bar{\nu}$, with $T\leq\frac{\mathrm{V}(\mu_{0})}{\alpha}$.
Proof.
Let $\mu_{0}\in\mathcal{P}_{2}(\mathbb{R}^{d})$. Thanks to Proposition 4.13, there exists $\mu\in\mathcal{A}_{[0,T_{0}]}(\mu_{0})$, with $T_{0}:=\displaystyle\frac{\mathrm{V}(\mu_{0})}{\alpha}$, such that $0\leq\mathrm{V}(\mu_{t})\leq-\alpha\,t+\mathrm{V}(\mu_{0})$, for all $t\in[0,T_{0}]$. In particular, the relation holds with $t=T_{0}$, leading to the equality $\mathrm{V}(\mu_{T_{0}})=0$. Hence, $\mu_{T_{0}}=\bar{\nu}$, by the assumptions on $\mathrm{V}$. ∎
@@ Appendix A A priori estimates
(A.1) $\begin{split}C_{[a,b]}(\bar{\mu})&:=e^{L(b-a)}\left(\mathsf{m}_{2}(\bar{\mu})+(K_{\mathrm{F}}(b-a)\right);\\ D_{[a,b]}(\bar{\mu})&:=\left(K_{\mathrm{F}}+L\,\left[e^{L(b-a)e^{L(b-a)}}C(\bar{\mu})+C(\bar{\mu})+L(b-a)e^{L(b-a)e^{L(b-a)}}C(\bar{\mu})\right]\right)^{2}.\end{split}$
@@ Appendix B Proof of Comparison Principle
In this section, devoted to the proof of Theorem 3.8, we recall some results proved in [JMQ20] which have been used in the proof of Propositions 4.8,4.13. We mention that the Comparison Principle in [JMQ20, Theorem 2] for (3.1) is proved in the context of bounded and uniformly continuous viscosity solutions $U:[0,T]\times\mathcal{P}_{2}(\mathbb{R}^{d})\to\mathbb{R}$. Due to the lack of boundedness in our framework, we readapt the result in Theorem 3.8.
@@
$-\xi:=\inf_{(t,\nu)\in[0,T]\times\mathcal{P}_{2}(\mathbb{R}^{d})}\left\{u_{2}(t,\nu)-u_{1}(t,\nu)\right\}$
and notice that by assumptions Theorem 3.8(1),(2) we have $-\xi\neq-\infty$. Assume by contradiction that $-\xi<0$ and fix $(t_{0},\nu_{0})\in X$ such that
@@
We start by showing that the first limit cannot be +$\infty$+ ∞. By contradiction, assume that there exist sequences $\{\varepsilon_{n}\}_{n},\,\{\delta_{n}\}_{n}$ with $\varepsilon_{n},\delta_{n}\to 0^{+}$ such that $\lim_{n\to+\infty}\rho_{\varepsilon_{n}\eta\delta_{n}}=+\infty$. Then, for $n$ sufficiently large we have $\varepsilon_{n},\delta_{n}\leq\frac{1}{2}$ and, by (B.4) and (B.5), recalling that $\eta<1$ we get
$-2R(L_{\bar{R},1}+L_{\bar{R},2})+m_{1}+m_{2}-(M_{1}+M_{2})+2\rho^{2}_{\varepsilon_{n}\eta\delta_{n}}<2\rho_{\varepsilon_{n}\eta\delta_{n}}.$
@@
$-2R(L_{\bar{R},1}+L_{\bar{R},2})+m_{1}+m_{2}-(M_{1}+M_{2})+\frac{\alpha^{2}}{\varepsilon_{n}}\leq 3\alpha(2\delta_{n}+\eta).$
Since the right hand side is bounded, while the left hand side tends to +$\infty$+ ∞, we get a contradiction. This proves the first relation in (B.6).
@@
By continuity of $u_{1},u_{2}$, the right hand side tends to 00 as $\varepsilon,\delta\to 0^{+}$, and this proves the second relation in (B.6).

### v2

@@ 5.1. The second moment as a Lyapunov function
Claim 1 is proved in Remark 3.9. In order to prove Claim 2, we proceed by proving that for any $\tilde{\mu}_{0}\in D$ there exists $\tilde{\mu}\in\mathcal{A}_{[0,+\infty)}(\mu_{0})$ such that the function $S\SS$ is nonincreasing, where
$D:=\left\{\nu=\sum_{i=1}^{N}\beta_{i}\,\delta_{x^{i}}\,:\,N\in\mathbb{N},\,(x^{i})_{i}\subset\mathbb{R}^{d},\,\beta_{i}\in[0,1],\,\sum_{i}\beta_{i}=1\right\}.$
Then, Claim 2 follows by density of $D$ into $\mathcal{P}_{2}(\mathbb{R}^{d})$ w.r.t. the metric $W_{2}$, relative compactness of $\bigcup_{\nu\in\mathcal{P}_{2}(\mathbb{R}^{d}),\mathsf{m}_{2}(\nu)\leq C}\mathcal{A}_{[0,T]}(\nu)$ for any $C,T>0$ and continuity of $\mathrm{V}$. Let $N\in\mathbb{N}$, $(\beta_{i})_{i=1,\dots,N}\subset[0,1]$ with $\sum_{i=1}^{N}\beta_{i}=1$ and $(x_{0}^{i})_{i=1,\dots,N}\subset\mathbb{R}^{d}$. Take $\tilde{\mu}_{0}=\sum_{i=1}^{N}\beta_{i}\delta_{x_{0}^{i}}$, then the trajectory $\tilde{\mu}=(\tilde{\mu}_{t})_{t}$, with
$\tilde{\mu}_{t}:=\sum_{i=1}^{N}\beta_{i}\delta_{e^{-\alpha t}x_{0}^{i}}$
is admissible for $\mathrm{F}$ starting from $\tilde{\mu}_{0}$. Moreover, for any $t\geq 0$ we have
$\mathrm{V}(\tilde{\mu}_{t})=\frac{1}{2}\mathsf{m}_{2}^{2}(\tilde{\mu}_{t})=\frac{1}{2}e^{-2\alpha t}\sum_{i=1}^{N}\beta_{i}|x_{0}^{i}|^{2}\leq\frac{1}{2}e^{-\alpha t}\mathsf{m}_{2}^{2}(\tilde{\mu}_{0})=e^{-\alpha t}\mathrm{V}(\tilde{\mu}_{0}).$
Following a similar reasoning as that reported at the end of the proof of Corollary 4.6, we conclude that $S\SS$, defined through $(\tilde{\mu}_{t})_{t}$, is nonincreasing in $0,+\infty)$.
We can give an intepretation of the proposed example e.g. in social sciences. Indeed, $\mathrm{F}$ can be interpreted as modeling individual behavioral adjustments in a social system, where $x\in\mathbb{R}^{d}$ represents the state of an agent (e.g., opinion, wealth, or position), and $\nu\in\mathcal{P}_{2}(\mathbb{R}^{d})$ denotes the distribution of the overall population. The available control at each state is given by a ball centered at the origin with radius proportional to both the agent’s deviation from consensus (represented by the origin) and the global dispersion of the population. The model captures situations where agents adjust their state under bounded effort and directional uncertainty, but with a responsiveness that increases with individual deviation and systemic disorder. The asymptotic reachability of the origin corresponds to a collective convergence toward consensus or centralization.
5.2. The squared Wasserstein distance as a Lyapunov function
This example, presented for simplicity in dimension $d=2$, is very similar to an example in [aver]. We consider the single-valued vector field $\mathrm{F}$ defined, for $x\in\mathbb{R}^{2}$ and $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$, as follows
$\mathrm{F}(x,\mu):=Ax+\int_{\mathbb{R}^{2}}By\,\mathrm{d}\mu(y),$
where $A:=\begin{pmatrix}0&1\\ -1&0\end{pmatrix}$ and $B$ a negative definite matrix such that for some $k>0$
$\langle Bz,z\rangle\leq-k|z|^{2}\quad\forall z\in\mathbb{R}^{2}.$
We will consider the associated continuity equation
(5.1) $\partial_{t}\mu_{t}+\mathrm{div}\left[\left(Ax+\int_{\mathbb{R}^{2}}By\,\mathrm{d}\mu_{t}(y)\right)\mu_{t}\right]=0\quad\mbox{for a.e. $t$ and $\mu_{t}$-a.e. $x\in\mathbb{R}^{2}$. }$
We consider $\bar{\nu}$ to be the probability measure with density $\frac{1}{c}e^{-(x_{1}^{2}+x^{2}_{2})}$ (where $c$ is a normalizing constant). It can be shown (cf. [aver, Section 4.1]) that $\bar{\nu}$ is an equilibria of (5.1). We wish to prove that $\mathrm{V}:\mathcal{P}_{2}(\mathbb{R}^{2})\to[0,+\infty)$, defined by $\mathrm{V}(\nu):=\frac{1}{2}W^{2}_{2}(\nu,\bar{\nu})$ is a Lyapunov function for $\mathrm{F}$ with rate $\varphi(r):=2kr$.
For this aim, we claim that $\mathrm{V}$ satisfies (4.1).
According to [Jimenez, Theorem 3.23, Proposition and Definition 3.1, Proposition 3.11, Remark 3.12], it is enough to show that
$\varphi(\mathrm{V}(\nu))+\mathcal{H}_{\mathrm{F}}(\nu,p)\leq 0,\qquad\forall\,\nu\in\mathcal{P}_{2}(\mathbb{R}^{d}),\,\forall\,p\in\partial_{\text{AG}}^{-}\mathrm{V}(\nu),$
where $\partial_{\text{AG}}^{-}\mathrm{V}(\nu)$ is defined in [ag] to be the set of all $p\in\mathrm{Tan}_{\nu}\mathcal{P}_{2}(\mathbb{R}^{2})$ (cf. [ags, (8.0.2)]) such that for all $\mu\in\mathcal{P}_{2}(\mathbb{R}^{2})$ and $\boldsymbol{\sigma}\in\Gamma_{o}(\nu,\mu)$,
$\mathrm{V}(\mu)-\mathrm{V}(\nu)\geq\int_{\mathbb{R}^{2}\times\mathbb{R}^{2}}\langle p(x),y-x\rangle\,\mathrm{d}\boldsymbol{\sigma}(x,y)+o(W_{2}(\mu,\nu)).$
Fix $\nu\in\mathcal{P}_{2}(\mathbb{R}^{d})$; combining [ags, Section 10.2, Theorem 10.2.6], [G2, Theorem 3.21, Corollary 3.22] and [Al-J, Theorem 3.2] it is possible to show that, if there exists a measurable function $T:\mathbb{R}^{d}\to\mathbb{R}^{d}$ such that $\Gamma_{o}(\nu,\bar{\nu})=\left\{(\mathrm{id}_{\mathbb{R}^{d}},T)_{\sharp}\nu\right\}$, then $\partial_{\text{AG}}^{-}\mathrm{V}(\nu)=\left\{\mathrm{id}_{\mathbb{R}^{d}}-T\right\}$; otherwise, $\partial_{\text{AG}}^{-}\mathrm{V}(\nu)=\emptyset$. Continuity of $\mathrm{V}$ ensures the existence of a dense subset of $\mathcal{P}_{2}(\mathbb{R}^{d})$ where $\partial_{\text{AG}}^{-}\mathrm{V}$ is not empty (cf. e.g. [JMQ21, Proposition 3.22]).
A straightforward adaptation of the argument in [aver, Section 4, Theorem 4.4] allows us to conclude that for all $p$ of the form $\mathrm{id}_{\mathbb{R}^{2}}-T$, with $T$ as above, we get
$2k\mathrm{V}(\nu)+\mathcal{H}_{\mathrm{F}}(\nu,p)\leq 0.$
Our claim is proved. By Theorem 4.7, $\bar{\nu}$ is asymptotically reachable for the dynamical system (5.1).
@@ Appendix A A priori estimates
(A.1) $\begin{split}C_{[a,b]}(\bar{\mu})&:=e^{L(b-a)}\left(\mathsf{m}_{2}(\bar{\mu})+(K_{\mathrm{F}}(b-a)\right);\\ D_{[a,b]}(\bar{\mu})&:=\left(K_{\mathrm{F}}+L\,\left[e^{L(b-a)e^{L(b-a)}}{\color[rgb]{0,0,0}C_{[a,b]}(\bar{\mu})}+{\color[rgb]{0,0,0}C_{[a,b]}(\bar{\mu})}+L(b-a)e^{L(b-a)e^{L(b-a)}}{\color[rgb]{0,0,0}C_{[a,b]}(\bar{\mu})}\right]\right)^{2}.\end{split}$
@@ Appendix B Proof of Comparison Principle
In this section, devoted to the proof of Theorem 3.11, we recall some results proved in [JMQ20] which have been used in the proof of Proposition 4.5. We mention that the Comparison Principle in [JMQ20, Theorem 2] for (3.1) is proved in the context of bounded and uniformly continuous viscosity solutions $U:[0,T]\times\mathcal{P}_{2}(\mathbb{R}^{d})\to\mathbb{R}$. Due to the lack of boundedness in our framework, we readapt the result in Theorem 3.11.
@@
$\xi:=\inf_{(t,\nu)\in[0,T]\times\mathcal{P}_{2}(\mathbb{R}^{d})}\left\{u_{2}(t,\nu)-u_{1}(t,\nu)\right\}.$
Assume by contradiction that $\xi<0$. Without loss of generality, we may assume that $\xi\neq-\infty$; otherwise, the proof proceeds analogously by replacing all occurrences of $\xi$ in what follows with an arbitrarily fixed $B>0$. Fix $(t_{0},\nu_{0})\in X$ such that
@@
We start by showing that the first limit cannot be $\infty$. By contradiction, assume that there exist sequences $\{\varepsilon_{n}\}_{n},\,\{\delta_{n}\}_{n}$ with $\varepsilon_{n},\delta_{n}\to 0^{+}$ such that $\lim_{n\to+\infty}\rho_{\varepsilon_{n}\eta\delta_{n}}=+\infty$. Then, for $n$ sufficiently large we have $\varepsilon_{n},\delta_{n}\leq\frac{1}{2}$ and, by (B.4) and (B.5), recalling that $\eta<1$ we get
$2R(L_{\bar{R},1}+L_{\bar{R},2})+m_{1}+m_{2}-(M_{1}+M_{2})+2\rho^{2}_{\varepsilon_{n}\eta\delta_{n}}<2\rho_{\varepsilon_{n}\eta\delta_{n}}.$
@@
$2R(L_{\bar{R},1}+L_{\bar{R},2})+m_{1}+m_{2}-(M_{1}+M_{2})+\frac{\alpha^{2}}{\varepsilon_{n}}\leq 3\alpha(2\delta_{n}+\eta).$
Since the right hand side is bounded, while the left hand side tends to $\infty$, we get a contradiction. This proves the first relation in (B.6).
@@
By continuity of $u_{1},u_{2}$, the right hand side tends to 0 as $\varepsilon,\delta\to 0^{+}$, and this proves the second relation in (B.6).